# Adapting Speech Foundation Models with Large Language Models for Unified Speech Recognition

### v1

@@ 4.3 AVSR results on noisy test sets
| ChatGLM v3-6B | iFLYTEK-Speech | 30.0 | 1.4 | 1.7 | 33.6 | 4.8 | |
@@
| Qwen 2.5-7B | iFLYTEK-Speech | 29.3 | 2.0 | 1.8 | 33.9 | 4.7 | |
@@
| ChatGLM v3-6B | iFLYTEK-Speech | 24.8 | 0.87 | 0.76 | 31.7 | 3.3 | |
@@
| Qwen 2.5-7B | iFLYTEK-Speech | 24.5 | 0.76 | 0.89 | 32.0 | 3.0 | |
@@
4.4 Generalization across SFMs and LLMs
To evaluate the generalization capability of our approach, we conducted experiments with different speech feature models (SFMs) and large language models (LLMs). Specifically, we replaced the WavLM-based SFM with two supervised learning alternatives: Whisper and iFLYTEK-Speech. Additionally, we compared ChatGLM v3-6B with Qwen 2.5-7B as the backbone LLM. For training, we used 433 hours of data for the visual injection pretraining stage and finetune the models on both 30 hours and 433 hours of the LRS3 dataset. We adopted a pretrained visual encoder (Base version) trained on 1759 hours of data. The experimental results are summarized in Table 3.
We first analyze the influence of different SFMs using the same LLM and identical amounts of finetuning data. The results show that iFLYTEK-Speech achieves the best ASR performance under both clean and noisy conditions, while WavLM performs best on the VSR task. For AVSR, WavLM achieves superior results on the noisy testset, whereas the clean AVSR results do not exhibit a clear pattern across different SFMs. When examining the influence of LLMs with the same SFM, we observe that ChatGLM v3-6B slightly outperforms Qwen 2.5-7B when using low-resource data (30 hours) for finetuning. However, this trend reverses with higher-resource data (433 hours), where Qwen 2.5-7B demonstrates better performance. Despite performance variations across different SFM and LLM combinations, our UASR-LLM consistently achieves significant improvements in recognition accuracy compared to baselines, demonstrating the strong generalization ability of our proposed method.
4.5 Ablation studies
| Labeled data | VIPT data | VSR | ASR | AVSR | ASR (N) | AVSR (N) |
@@ 4.5 Ablation studies
4.5.1 Visual injection pretraining
Table 4 summarizes the experimental results of our visual injection pretraining ablation study. The results demonstrate that omitting the proposed visual injection pretraining stage (0h for VIPT data) leads to degraded performance in both VSR and AVSR tasks. Furthermore, increasing the amount of pretraining data during the visual injection pretraining stage yields additional improvements in VSR and AVSR performance. These superior results for VSR and AVSR tasks validate the effectiveness of our pretraining stage in processing visual information. Interestingly, for the ASR task, the model without the pretraining stage achieves the best results. Since we froze all SFM parameters throughout training, the audio processing capability of SFM itself remained unaffected by the visual injection pretraining stage. This phenomenon may be attributed to our parameter sharing strategy for USR in LLMs, where enhanced AVSR and VSR performance appears to cause a slight degradation in ASR performance—a trade-off we explore in the following section.
| Labeled data | Method | VSR | ASR | AVSR | ASR (N) | AVSR (N) |
@@ 4.5.2 Parameter sharing for USR
This section compares our parameter sharing approach for realizing USR in a single model with the traditional method of training separate models for each task. The results (w/o shared parameters) are presented in Table 5. The table shows that training separate models leads to better ASR performance but worse VSR performance. Overall, the proposed USR method achieves comparable performance to the traditional approach that optimizes individual models for each task. These results demonstrate the effectiveness of our method for realizing USR, significantly reducing both training and deployment resources while maintaining strong performance across all tasks.
4.5.3 Integration of the LLM
@@
This section compares our proposed method based on a decoder-only LLM with the traditional approach using a 6-layer Transformer-based decoder with cross-attention. The results (w/o LLM) are shown in Table 5. The table demonstrates that our proposed method employing LLMs significantly outperforms the traditional Transformer decoder. This validates that the strong contextual and language modeling capabilities of LLMs, learned from large text corpora, can be effectively transferred to enhance speech recognition performance, which has also been demonstrated in previous studies [36].
5 Conclusion
@@
In this work, we propose UASR-LLM, a novel approach that adapts speech foundation models (SFMs) for unified speech recognition, encompassing ASR, VSR, and AVSR within a single framework using a decoder-only LLM for text prediction. Our method augments the SFM with a visual encoder and inserts visual injection modules into each SFM block to fuse visual representations with audio features. We introduce a visual injection pretraining stage where corrupted audio-visual inputs are processed to predict clean audio representations, followed by speech recognition finetuning with modality dropout. The approach leverages the complementary strengths of SFMs pretrained on large-scale audio corpora and LLMs pretrained on extensive text data.
Experimental results demonstrated the effectiveness of UASR-LLM, achieving superior performance compared to state-of-the-art baselines across ASR, VSR, and AVSR tasks under both clean and noisy conditions. Our method significantly improved recognition robustness across varying signal-to-noise ratios and noise types. Experiments with different SFM and LLM combinations show consistent improvements over baselines, demonstrating the generalization capability of our framework. Ablation studies validated the importance of visual injection pretraining, parameter sharing, and LLM integration.
Despite strong performance, our method has several limitations that warrant future investigation. First, while LLM integration leads to substantial speech recognition improvements, it also requires significant computational resources and memory footprint. This computational burden may limit practical deployment, suggesting the need for more efficient approaches such as dynamic pooling of adjacent audio-visual representations to reduce input tokens or exploring lightweight LLM architectures. Second, our method still lags behind previous lipreading studies that employ much larger audio-visual datasets, indicating that the combination of SFMs and LLMs cannot fully compensate for the shortage of large-scale audio-visual training corpora. This limitation fundamentally affects the visual representation extraction capacity of the visual encoder. We anticipate that employing visual encoders pretrained on large-scale visual datasets could further improve our results. Future work will focus on addressing these challenges to enhance both computational efficiency and performance of unified speech recognition systems.
6 Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant No. 62401348, 62206162), Fundamental Research Funds for the Central Universities (Grant No. GK202406005), and Young Talent Fund of Association for Science and Technology in Shaanxi, China (Grant No. 20250126).
7 Declaration of Generative AI and AI-assisted technologies in the writing process
@@
During the preparation of this work, the authors used iFLYTEK’s Xinghuo777https://xinghuo.xfyun.cn/ in order to improve readability and language. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.

### v2

@@ IV-C AVSR results on noisy test sets
| ChatGLM v3-6B | Conformer-Speech | 30.0 | 1.4 | 1.7 | 33.6 | 4.8 | |
@@
| Qwen 2.5-7B | Conformer-Speech | 29.3 | 2.0 | 1.8 | 33.9 | 4.7 | |
@@
| ChatGLM v3-6B | Conformer-Speech | 24.8 | 0.87 | 0.76 | 31.7 | 3.3 | |
@@
| Qwen 2.5-7B | Conformer-Speech | 24.5 | 0.76 | 0.89 | 32.0 | 3.0 | |
@@
IV-D Generalization across SFMs and LLMs
To evaluate the generalization capability of our approach, we conducted experiments with different speech feature models (SFMs) and large language models (LLMs). Specifically, we replaced the WavLM-based SFM with two supervised learning alternatives: Whisper and Conformer-Speech. Additionally, we compared ChatGLM v3-6B with Qwen 2.5-7B as the backbone LLM. For training, we used 433 hours of data for the visual injection pretraining stage and finetune the models on both 30 hours and 433 hours of the LRS3 dataset. We adopted a pretrained visual encoder (Base version) trained on 1759 hours of data. The experimental results are summarized in Table III.
We first analyze the influence of different SFMs using the same LLM and identical amounts of finetuning data. The results show that Conformer-Speech achieves the best ASR performance under both clean and noisy conditions, while WavLM performs best on the VSR task. For AVSR, WavLM achieves superior results on the noisy testset, whereas the clean AVSR results do not exhibit a clear pattern across different SFMs. When examining the influence of LLMs with the same SFM, we observe that ChatGLM v3-6B slightly outperforms Qwen 2.5-7B when using low-resource data (30 hours) for finetuning. However, this trend reverses with higher-resource data (433 hours), where Qwen 2.5-7B demonstrates better performance. Despite performance variations across different SFM and LLM combinations, our UASR-LLM consistently achieves significant improvements in recognition accuracy compared to baselines, demonstrating the strong generalization ability of our proposed method.
IV-E Ablation studies
| Labeled data | VIPT data | VSR | ASR | AVSR | ASR (Noisy) | AVSR (Noisy) |
@@ IV-E Ablation studies
IV-E1 Visual injection pretraining
Table IV summarizes the experimental results of our visual injection pretraining ablation study. The results demonstrate that omitting the proposed visual injection pretraining stage (0h for VIPT data) leads to degraded performance in both VSR and AVSR tasks. Furthermore, increasing the amount of pretraining data during the visual injection pretraining stage yields additional improvements in VSR and AVSR performance. These superior results for VSR and AVSR tasks validate the effectiveness of our pretraining stage in processing visual information. Interestingly, for the ASR task, the model without the pretraining stage achieves the best results. Since we froze all SFM parameters throughout training, the audio processing capability of SFM itself remained unaffected by the visual injection pretraining stage. This phenomenon may be attributed to our parameter sharing strategy for USR in LLMs, where enhanced AVSR and VSR performance appears to cause a slight degradation in ASR performance—a trade-off we explore in the following section.
| Labeled data | Method | VSR | ASR | AVSR | ASR (Noisy) | AVSR (Noisy) |
@@ IV-E2 Parameter sharing for USR
This section compares our parameter sharing approach for realizing USR in a single model with the traditional method of training separate models for each task. The results (w/o shared parameters) are presented in Table V. The table shows that training separate models leads to better ASR performance but worse VSR performance. Overall, the proposed USR method achieves comparable performance to the traditional approach that optimizes individual models for each task. These results demonstrate the effectiveness of our method for realizing USR, significantly reducing both training and deployment resources while maintaining strong performance across all tasks.
IV-E3 Integration of the LLM
@@
This section compares our proposed method based on a decoder-only LLM with the traditional approach using a 6-layer Transformer-based decoder with cross-attention. The results (w/o LLM) are shown in Table V. The table demonstrates that our proposed method employing LLMs significantly outperforms the traditional Transformer decoder. This validates that the strong contextual and language modeling capabilities of LLMs, learned from large text corpora, can be effectively transferred to enhance speech recognition performance, which has also been demonstrated in previous studies [46].
V Conclusion
@@
In this work, we propose UASR-LLM, a novel approach that adapts speech foundation models (SFMs) for unified multimodal speech recognition, encompassing ASR, VSR, and AVSR within a single framework using a decoder-only LLM for text prediction. Our method augments the SFM with a visual encoder and inserts visual injection modules into each SFM block to fuse visual representations with audio features. We introduce a visual injection pretraining stage where corrupted audio-visual inputs are processed to predict clean audio representations, followed by speech recognition finetuning with modality dropout. The approach leverages the complementary strengths of SFMs pretrained on large-scale audio corpora and LLMs pretrained on extensive text data. Experimental results demonstrated the effectiveness of UASR-LLM, achieving superior performance compared to state-of-the-art baselines across ASR, VSR, and AVSR tasks under both clean and noisy conditions. Our method significantly improved recognition robustness across varying signal-to-noise ratios and noise types. Experiments with different SFM and LLM combinations show consistent improvements over baselines, demonstrating the generalization capability of our framework. Ablation studies validated the importance of visual injection pretraining, parameter sharing, and LLM integration. In the future study, we plan to focus on developing more efficient approaches such as dynamic pooling of adjacent audio-visual representations to reduce input tokens or exploring lightweight LLM architectures.
VI Acknowledgements
@@
During the preparation of this manuscript, the authors used iFLYTEK’s Xinghuo777https://xinghuo.xfyun.cn for language editing and readability improvement. After using this tool, the authors carefully reviewed and revised the content as necessary and take full responsibility for the accuracy and integrity of the final manuscript.